OpenStudy (anonymous):

Simplify using your knowledge of properties. [(1/3)(3)](2 + -1 + 1)

OpenStudy (anonymous):

first is the multiplicative inverse, (1/3)*3=1 second is the additive inverse, -1+1=0 1(2)=2

jhonyy9 (jhonyy9):

(1/3)(3)=3/3 =1(2+(-1+1))=1(2+0)=2